If the perimiter of a rectangle with a long side of 2x+1 and a shorter side of x-3 is 44 then x = ....
Answer by nerdybill(7384)

If the perimiter of a rectangle with a long side of 2x+1 and a shorter side of x-3 is 44 then x = ....
.
Since:
perimeter = 2(width + length)
44 = 2((x-3) + (2x+1))
44 = 2(x-3 + 2x+1)
44 = 2(3x-2)
22 = 3x-2
24 = 3x
8 = x